# What Should I Cite? A RAG Benchmark for Academic Citation Prediction

### v1

@@ Abstract.
With the rapid growth of Web-based academic publications, more and more papers are being published annually, making it increasingly difficult to find relevant prior work. Citation prediction aims to automatically suggest appropriate references, helping scholars navigate the expanding scientific literature. Here we present CiteRAG, the first comprehensive retrieval-augmented generation (RAG)-integrated benchmark for evaluating large language models on academic citation prediction, featuring a multi-level retrieval strategy, specialized retrievers, and generators. Our benchmark makes four core contributions: (1) We establish two instances of the citation prediction task with different granularity. Task 1 focuses on coarse-grained list-specific citation prediction, while Task 2 targets fine-grained position-specific citation prediction. To enhance these two tasks, we build a dataset containing 7,267 instances for Task 1 and 8,541 instances for Task 2, enabling comprehensive evaluation of both retrieval and generation. (2) We construct a three-level large-scale corpus with 554k papers spanning many major subfields, using an incremental pipeline. (3) We propose a multi-level hybrid RAG approach for citation prediction, fine-tuning embedding models with contrastive learning to capture complex citation relationships, paired with specialized generation models. (4) We conduct extensive experiments across state-of-the-art language models, including closed-source APIs, open-source models, and our fine-tuned generators, demonstrating the effectiveness of our framework. Our open-source toolkit enables reproducible evaluation and focuses on academic literature, providing the first comprehensive evaluation framework for citation prediction and serving as a methodological template for other scientific domains. Our source code and data are released at https://github.com/LQgdwind/CiteRAG.
@@ 1. Introduction
Despite extensive research on citation prediction (Cohan et al., 2019; Zhang et al., 2023), there remains a notable absence of a comprehensive benchmark for evaluating models in this domain. Existing benchmarks suffer from several key limitations: (1) oversimplified task formulations, as real-world users often query related studies using specific sections of a paper, such as the introduction, methodology, or experimental details, to identify the most relevant and fine-grained prior work (Cohan et al., 2019), whereas current benchmarks typically treat citation prediction as a binary classification or document-level retrieval problem within a fixed corpus, an oversimplification that diverges from practical scenarios; (2) ignorance of the hierarchical structure of scientific literature, since scientific texts are inherently layered and rhetorically organized (Teufel et al., 1999; Zhang et al., 2023; Kashyap et al., 2023), yet current benchmarks often flatten this complexity into a textual knowledge graph assuming full-text availability, thus neglecting the multi-level organization of scientific discourse; and (3) suboptimal methodological design, as most existing approaches rely on fine-tuning simple embedding-based retrievers (Reimers and Gurevych, 2019; Gao et al., 2021; Lee et al., 2019), which struggle to capture the nuanced, structural nature of citation relationships and tend to lose retrieval effectiveness as the candidate set expands (Lee et al., 2019); (4) a lack of systematic evaluation framework, as current benchmarks typically focus on citation network construction and simple embedding-based methods while failing to provide a comprehensive evaluation of more advanced approaches such as retrieval-augmented generation (RAG) (Lewis et al., 2020; Izacard and Grave, 2021). Other platforms, such as Deep Research-style agents (Zheng et al., 2025d), focus on literature exploration rather than serving as standardized benchmarks, and cannot serve as formal benchmarks for citation prediction.
@@
To address this gap, we introduce a comprehensive benchmark for citation prediction that encompasses two complementary tasks operating at different levels of granularity. List-specific citation prediction (Task 1) evaluates a system’s ability to generate complete reference lists for papers. Position-specific citation prediction (Task 2) assesses a system’s capability to predict precise citations for individual reference placeholders within the paper text. Building on these two tasks, we collected a large corpus of scientific texts from Google Scholar, resulting in a multidisciplinary, hierarchical corpus of 554k papers. Each paper is represented with multi-level textual information and associated with a large candidate set. To facilitate the evaluation of citation prediction tasks, we further curated 7,267 samples for Task 1 and 8,541 samples for Task 2.
Given the massive size of the candidate pool and the limitations of conventional embedding-based methods in capturing citation relations, we design a multi-level hybrid retrieval-augmented generation (RAG) approach for citation prediction. Our approach integrates a carefully fine-tuned, state-of-the-art 8B retriever with a suite of task-specific large language models (LLMs). To establish a solid foundation for future research, we conduct a comprehensive evaluation of our RAG pipeline alongside a wide range of baseline methods, including traditional embedding-based retrievers and both open- and closed-source LLMs. The results highlight the intrinsic complexity of the citation prediction task and confirm the effectiveness of our proposed approach. Finally, we release an open-source toolkit that enables reproducible evaluation and provides the first comprehensive framework dedicated to citation prediction. Table 1 summarizes the key features of our benchmark compared to existing datasets. Our benchmark is multidisciplinary and supports RAG-based baselines. It also includes the evaluation for hallucination and diversity, is structured as a hierarchical RAG pipeline, and provides a reproducible open-source toolkit.
@@
(1) Task Definition and Granularity Distinction. We define two novel instances of the citation prediction task: Task 1, focusing on coarse-grained list-specific prediction, and Task 2, targeting fine-grained position-specific prediction. The dataset we built contains 7,267 instances for Task 1 and 8,541 instances for Task 2, enabling comprehensive evaluation of both retrieval and generation across different task granularities. (Section 2 and Section 3)
(2) Multi-Level Incremental Corpus Construction. We construct a large-scale hierarchical corpus with 554,719 papers, built using an incremental pipeline that supports continuous updates. This enables flexible access to multi-level information and scalable expansion of the academic knowledge base for citation prediction tasks. (Section 3)
(3) Multi-Level Hybrid Retrieval and Generation Framework. We propose a multi-level hybrid RAG approach, combining specialized retrievers tuned for different corpus levels with generation models fine-tuned to capture complex citation relationships. This multi-level hybrid RAG approach improves both retrieval precision and context-aware citation generation compared to single-level methods. (Section 4)
(4) Open-Source RAG Evaluation Toolkit. We establish a comprehensive evaluation framework covering the full citation prediction pipeline, from retrieval effectiveness to end-to-end generation performance. The toolkit provides standardized metrics, evaluation protocols, and extensive comparison across state-of-the-art models, including closed-source APIs, open-source LLMs, and our fine-tuned generators, enabling reproducible benchmarking and facilitating further research. We hope this toolkit serves as a foundation for future work. (Section 5, Section 6 and Section 7)
@@ 2.2. Task 2: Position-Specific Citation Prediction
This fine-grained formulation requires understanding the immediate textual context surrounding each reference placeholder. The prediction system must analyze specific claims, methodologies, or concepts being discussed, identify the required citation support type, and select the most appropriate reference. This reflects the actual writing process where authors make precise decisions about citations at each point in their argumentation, considering chronological precedence, methodological relevance, and argumentative support.
@@ 3.1. Data Collection
To construct a comprehensive corpus for academic citation prediction, we systematically collected papers from Google Scholar over the past decade, ensuring broad temporal coverage of evolving research trends and citation patterns. The collection framework implements adaptive query strategies using year-based filtering for categories with fewer than 2,000 papers, month-by-month querying for 2,000-12,000 papers, and week-by-week collection for categories exceeding 12,000 papers to prevent API-related data loss.
@@ 3.2.2. Dataset Pre-processing
The Position-Specific Citation Prediction (Task 2) preprocessing employs a more sophisticated multi-stage approach to create fine-grained citation scenarios. We also initially filter out papers with non-standard citations, then analyze citation distribution patterns to identify the three most frequently cited sections within each paper. For these high-citation sections, we preserve the top three most frequent citations with valid identifiers while removing other citation markers, creating clean textual contexts with precisely positioned reference placeholders. Papers where any single reference appears more than ten times within a section are excluded to prevent citation bias.
@@ 4.1. Multi-Level Hybrid Retrieval Strategy
Our chunking strategy employs title and abstract content as standardized query input, maintaining consistency with Task 1 requirements while enabling efficient retrieval across all corpus granularities. Given query paper consisting of title and abstract, the retrieval system performs parallel top-k similarity search across the three pre-established corpus levels:
@@ 4.2. Contrastive Learning Fine-tuning
Standard embedding models are inadequately optimized for academic citation relationships, which encompass complex interdependencies beyond semantic similarity. We fine-tune Qwen-3-embedding-8B as CitationRetriever-8B using contrastive learning to capture citation-specific patterns.
@@ 5.1. Retriever Evaluation Metrics
We employ two standard information retrieval metrics to evaluate retriever performance: Recall@k and Mean Reciprocal Rank@k (MRR@k). Detailed mathematical formulations are provided in Appendix B.
@@ 5.2. Task 1 Evaluation Metrics
Task 1 evaluation employs three standard metrics for reference generation: Recall@k, Normalized Discounted Cumulative Gain@k (NDCG@k), and Hit@k. Complete definitions are provided in Appendix B.
@@ 5.3. Task 2 Evaluation Metrics
where is the total number of reference placeholders, correct@k denotes correct predictions within top-k, and is the position of correct prediction .
@@ 8.1. Retrieval-Augmented Generation
Traditional retrieval methods (Zheng et al., 2025b, a, c) are no longer sufficient to meet the demands of the era of large language models (Li et al., 2025; Feng et al., 2025; Zhang et al., 2025b; Team, 2025; Hui et al., 2024; Zhang et al., 2026, 2025a; Kang et al., 2026; He et al., 2026), retrieval-augmented generation (RAG) has become a key paradigm for mitigating large language models’ limitations in domain-specific or time-sensitive tasks, where models often hallucinate or generate outdated responses (Tonmoy et al., 2024). RAG addresses this via a three-stage pipeline: retrieval from external knowledge bases, query–context integration, and grounded response generation (Asai et al., 2024). Subsequent work improves RAG through enhanced retrievers (Lewis et al., 2020), adaptive retrieval strategies (Jeong et al., 2024), reasoning-based prompting such as IR-CoT (Trivedi et al., 2023), and graph-structured retrieval like GraphRAG (Edge et al., 2024). However, these general frameworks lack the hierarchical corpus organization and multi-granularity retrieval needed for academic literature, where citation reasoning spans titles, abstracts, and full texts.
@@ 8.2. Citation Generation and Evaluation
Recent RAG systems have also enabled citation-aware text generation for verifiability. Early models such as WebGPT (Nakano et al., 2021) and reinforcement-based methods (Menick et al., 2022) laid the groundwork, while later systems including ReGen (Qian et al., 2023) and WebCPM (Qin et al., 2023) refined citation integration through improved training. Recent advances in few-shot learning, self-reflection mechanisms, and fine-grained reward modeling have further improved citation quality. Evaluation has evolved from manual annotation (AIS (Rashkin et al., 2023)) to automated and benchmark-based approaches such as ALCE (Gao et al., 2023). However, existing work mainly targets Web-based question answering and lacks support for academic citation prediction, which requires both document-level reference modeling and position-specific citation reasoning.
@@ Appendix B Detailed Metric Definitions
For the retriever evaluation, Recall@k measures the proportion of relevant papers successfully retrieved within the top-k positions:
@@
For the task 1 evaluation, Hit@k quantifies the absolute number of correct predictions within the top-k results.
@@ C.3. Noise Robustness Analysis
Table 4 evaluates model robustness against retrieval noise by introducing varying proportions of irrelevant papers into retrieved contexts. Closed-source models including GPT-5, Claude-4-Sonnet, and Gemini-2.5-Pro demonstrate moderate robustness, maintaining 80-90% of their clean performance at 20% noise levels but declining to 75-85% at higher noise ratios. CitationGenerator-30B demonstrates superior robustness due to its fine-tuning, retaining approximately 80% of clean performance even at 40% noise and maintaining reasonable effectiveness at 80% noise levels where other models collapse. These results demonstrate that retrieval noise induces gradual performance degradation rather than complete system failure, validating the robustness of RAG-based citation prediction in realistic noisy retrieval environments.
@@ C.5. Difficult Case Analysis
Figure 8 illustrates two challenging scenarios that expose current system limitations. The left panel shows a complex nested clause case where the citation appears within a deeply embedded subordinate clause, causing models to focus on the main clause about generalization rather than the nested context discussing momentum-based updates, resulting in incorrect predictions. The right panel demonstrates technical detail overload, where multiple scattered concepts in a lengthy sentence distract models from identifying the correct transformer mechanism reference, leading to recency bias toward superficial keyword matches.

### v2

@@ Abstract.
With the rapid growth of Web-based academic publications, more and more papers are being published annually, making it increasingly difficult to find relevant prior work. Citation prediction aims to automatically suggest appropriate references, helping scholars navigate the expanding scientific literature. Here we present CiteRAG, the first comprehensive retrieval-augmented generation (RAG)-integrated benchmark for evaluating large language models on academic citation prediction, featuring a multi-level retrieval strategy, specialized retrievers, and generators. Our benchmark makes four core contributions: (1) We establish two instances of the citation prediction task with different granularity. Task 1 focuses on coarse-grained list-specific citation prediction, while Task 2 targets fine-grained position-specific citation prediction. To enhance these two tasks, we build a dataset containing 7,267 instances for Task 1 and 8,541 instances for Task 2, enabling comprehensive evaluation of both retrieval and generation. (2) We construct a three-level large-scale corpus with 554k papers spanning many major subfields, using an incremental pipeline. (3) We propose a multi-level hybrid RAG approach to citation prediction, fine-tuning embedding models with contrastive learning to capture complex citation relationships, paired with specialized generation models. (4) We conduct extensive experiments across state-of-the-art language models, including closed-source APIs, open-source models, and our fine-tuned generators, demonstrating the effectiveness of our framework. Our open-source toolkit enables reproducible evaluation and focuses on academic literature, providing the first comprehensive evaluation framework for citation prediction and serving as a methodological template for other scientific domains. Our source code and data are released at https://github.com/LQgdwind/CiteRAG.
@@ 1. Introduction
Despite extensive research on citation prediction (Cohan et al., 2019; Zhang et al., 2023), there remains a notable absence of a comprehensive benchmark for evaluating models in this domain. Existing benchmarks suffer from several key limitations: (1) oversimplified task formulations, as real-world users often query related studies using specific sections of a paper, such as the introduction, methodology, or experimental details, to identify the most relevant and fine-grained prior work (Cohan et al., 2019), whereas current benchmarks typically treat citation prediction as a binary classification or document-level retrieval problem within a fixed corpus, an oversimplification that diverges from practical scenarios; (2) ignorance of the hierarchical structure of scientific literature, since scientific texts are inherently layered and rhetorically organized (Teufel et al., 1999; Zhang et al., 2023; Kashyap et al., 2023), yet current benchmarks often flatten this complexity into a textual knowledge graph assuming full-text availability, thus neglecting the multi-level organization of scientific discourse; (3) suboptimal methodological design, as most existing approaches rely on fine-tuning simple embedding-based retrievers (Reimers and Gurevych, 2019; Gao et al., 2021; Lee et al., 2019), which struggle to capture the nuanced, structural nature of citation relationships and tend to lose retrieval effectiveness as the candidate set expands (Lee et al., 2019); (4) a lack of systematic evaluation framework, as current benchmarks typically focus on citation network construction and simple embedding-based methods while failing to provide a comprehensive evaluation of more advanced approaches such as retrieval-augmented generation (RAG) (Lewis et al., 2020; Izacard and Grave, 2021). Other platforms, such as Deep Research-style agents (Zheng et al., 2025d), focus on literature exploration rather than serving as standardized benchmarks, and cannot serve as formal benchmarks for citation prediction.
@@
To address this gap, we propose a comprehensive benchmark for citation prediction that encompasses two complementary tasks operating at different levels of granularity. List-specific citation prediction (Task 1) evaluates a system’s ability to generate complete reference lists for papers. Position-specific citation prediction (Task 2) assesses a system’s capability to predict precise citations for individual reference placeholders within the paper text. Building on these two tasks, we collected a large corpus of scientific texts from Google Scholar, resulting in a multidisciplinary, hierarchical corpus of 554k papers. Each paper is represented with multi-level textual information and associated with a large candidate set. To facilitate the evaluation of citation prediction tasks, we further curated 7,267 samples for Task 1 and 8,541 samples for Task 2.
Given the massive size of the candidate pool and the limitations of conventional embedding-based methods in capturing citation relations, we design a multi-level hybrid retrieval-augmented generation (RAG) approach to citation prediction. Our approach integrates a carefully fine-tuned, state-of-the-art 8B retriever with a suite of task-specific large language models (LLMs). To establish a solid foundation for future research, we conduct a comprehensive evaluation of our RAG pipeline alongside a wide range of baseline methods, including both open- and closed-source LLMs as well as traditional embedding-based retrievers. The results highlight the intrinsic complexity of the citation prediction tasks and confirm the effectiveness of our proposed approach. Finally, we release an open-source toolkit that enables reproducible evaluation and provides the first comprehensive framework dedicated to citation prediction. Table 1 summarizes the key features of our benchmark compared to existing datasets. Our benchmark is multidisciplinary and supports RAG-based baselines. It also includes the evaluation for hallucination and diversity, is structured as a hierarchical RAG pipeline, and provides a reproducible open-source toolkit.
@@
(1) Task Definition and Granularity Distinction. We define two novel instances of the citation prediction tasks: Task 1, focusing on coarse-grained list-specific prediction, and Task 2, targeting fine-grained position-specific prediction. The dataset we built contains 7,267 instances for Task 1 and 8,541 instances for Task 2, enabling comprehensive evaluation of both retrieval and generation across different task granularities (Section 2 and Section 3).
(2) Multi-Level Incremental Corpus Construction. We construct a large-scale hierarchical corpus with 554,719 papers, built using an incremental pipeline that supports continuous updates. This enables flexible access to multi-level information and scalable expansion of the academic knowledge base for citation prediction tasks (Section 3).
(3) Multi-Level Hybrid Retrieval and Generation Framework. We propose a multi-level hybrid RAG approach, combining specialized retrievers tuned for different corpus levels with generation models fine-tuned to capture complex citation relationships. This multi-level hybrid RAG approach improves both retrieval precision and context-aware citation generation compared to single-level methods (Section 4).
(4) Open-Source RAG Evaluation Toolkit. We establish a comprehensive evaluation framework covering the full citation prediction pipeline, from retrieval effectiveness to end-to-end generation performance. The toolkit provides standardized metrics, evaluation protocols, and extensive comparison across state-of-the-art models, including closed-source APIs, open-source LLMs, and our fine-tuned generators, enabling reproducible benchmarking and facilitating further research (Section 5, Section 6 and Section 7).
@@ 2.2. Task 2: Position-Specific Citation Prediction
This fine-grained formulation requires understanding the immediate textual context surrounding each reference placeholder. The prediction system must analyze specific claims, methodologies, or concepts being discussed, identify the required citation support type, and select the most appropriate reference. This prediction system reflects the actual writing process where authors make precise decisions about citations at each point in their argumentation, considering chronological precedence, methodological relevance, and argumentative support.
@@ 3.1. Data Collection
To construct a comprehensive corpus for academic citation prediction, we systematically collected computer science papers from Google Scholar over the past decade, ensuring broad temporal coverage of evolving research trends and citation patterns.
The collection framework implements dynamic query strategies that adapt to the volume characteristics of different research areas. For smaller subcategories containing fewer than 2,000 papers, direct queries with year-based filtering provide efficient retrieval. Medium-sized categories with 2,000 to 12,000 papers employ month-by-month querying. Large categories exceeding 12,000 papers utilize week-by-week collection strategies to prevent data loss due to API limitations. This adaptive approach successfully collected papers across 10 computer science subcategories, including AI & ML, DB & DM, Systems, and others.
@@ 3.2.2. Dataset Pre-processing
The Position-Specific Citation Prediction (Task 2) preprocessing employs a more sophisticated multi-stage approach to create fine-grained citation scenarios. We also initially filter out papers with non-standard citations, and then analyze citation distribution patterns to identify the three most frequently cited sections within each paper. For these high-citation sections, we preserve the top three most frequent citations with valid identifiers while removing other citation markers, creating clean textual contexts with precisely positioned reference placeholders. Papers where any single reference appears more than ten times within a section are excluded to prevent citation bias.
@@ 4.1. Multi-Level Hybrid Retrieval Strategy
Our chunking strategy employs title and abstract content as standardized query input, maintaining consistency with Task 1 requirements while enabling efficient retrieval across all corpus granularities. Given query paper consisting of title and abstract, the retrieval system performs parallel top- similarity search across the three pre-established corpus levels:
@@ 4.2. Contrastive Learning Fine-tuning
Standard embedding models are inadequately optimized for academic citation relationships, which encompass complex interdependencies beyond semantic similarity. Then, we fine-tune Qwen-3-embedding-8B as CitationRetriever-8B using contrastive learning to capture citation-specific patterns.
@@ 5.1. Retriever Evaluation Metrics
We employ two standard information retrieval metrics to evaluate retriever performance: Recall@k and Mean Reciprocal Rank@k (MRR@k). The detailed mathematical formulations are provided in Appendix B.
@@ 5.2. Task 1 Evaluation Metrics
Task 1 evaluation employs three standard metrics for reference generation: Recall@k, Normalized Discounted Cumulative Gain@k (NDCG@k), and Hit@k. The complete definitions are provided in Appendix B.
@@ 5.3. Task 2 Evaluation Metrics
where is the total number of reference placeholders, correct@k denotes correct predictions within top-, and is the position of correct prediction .
@@ 8.1. Retrieval-Augmented Generation
Traditional retrieval methods (Zheng et al., 2025b, a, c) are no longer sufficient to meet the demands of the era of large language models (Li et al., 2025; Feng et al., 2025; Zhang et al., 2025b; Team, 2025; Hui et al., 2024; Zhang et al., 2026b, 2025a; Kang et al., 2026; He et al., 2026; Zhang et al., 2026a; Wang et al., 2025), retrieval-augmented generation (RAG) has become a key paradigm for mitigating large language models’ limitations in domain-specific or time-sensitive tasks, where models often hallucinate or generate outdated responses (Tonmoy et al., 2024). RAG addresses these limitations via a three-stage pipeline: retrieval from external knowledge bases, query–context integration, and grounded response generation (Asai et al., 2024). Subsequent work improves RAG through enhanced retrievers (Lewis et al., 2020), adaptive retrieval strategies (Jeong et al., 2024), reasoning-based prompting such as IR-CoT (Trivedi et al., 2023), and graph-structured retrieval like GraphRAG (Edge et al., 2024). However, these general frameworks lack the hierarchical corpus organization and multi-granularity retrieval needed for academic literature, where citation reasoning spans titles, abstracts, and full texts.
@@ 8.2. Citation Generation and Evaluation
Recent RAG systems have enabled citation-aware text generation for verifiability. Early models such as WebGPT (Nakano et al., 2021) and RL-based methods (Menick et al., 2022) laid the groundwork, while later systems including ReGen (Qian et al., 2023) and WebCPM (Qin et al., 2023) refined citation integration through improved training. Recent advances in few-shot learning, self-reflection mechanisms, and fine-grained reward modeling have further improved citation quality. Evaluation has evolved from manual annotation (AIS (Rashkin et al., 2023)) to automated and benchmark-based approaches such as ALCE (Gao et al., 2023). However, existing work mainly targets Web-based question answering and lacks support for academic citation prediction, which requires both document-level reference modeling and position-specific citation reasoning.
@@ Appendix B Detailed Metric Definitions
For the retriever evaluation, Recall@k measures the proportion of relevant papers successfully retrieved within the top- positions:
@@
For the task 1 evaluation, Hit@k quantifies the absolute number of correct predictions within the top- results.
@@ C.3. Noise Robustness Analysis
Table 4 evaluates model robustness against retrieval noise by introducing varying proportions of irrelevant papers into retrieved contexts. Closed-source models including GPT-5, Claude-4-Sonnet, and Gemini-2.5-Pro demonstrate moderate robustness, maintaining 80-90% of their clean performance at 20% noise levels but declining to 75-85% at higher noise ratios. CitationGenerator-30B demonstrates superior robustness due to its fine-tuning, retaining approximately 80% of clean performance even at 40% noise and maintaining reasonable effectiveness at 80% noise levels where other models collapse.
@@ C.5. Difficult Case Analysis
Figure 8 illustrates two challenging scenarios that expose current system limitations. The left panel shows a complex nested clause case where the citation appears within a deeply embedded subordinate clause, causing models to focus on the main clause about generalization rather than the nested context discussing momentum-based updates, resulting in incorrect predictions. The right panel demonstrates technical detail overload, where multiple scattered concepts in a lengthy sentence distract models from identifying the correct transformer mechanism reference, leading to recency bias toward superficial keyword matches. These failure cases highlight critical challenges in handling syntactically complex contexts and filtering relevant information from dense technical descriptions.